Journalist Report – January 6th

Sol 8 – 01/06/2018

Name the space movie (or show) given the following quote. Answer at the end of the Report:

“Space,” it says, “is big. Really big. You just won’t believe how vastly, hugely, mindbogglingly big it is. I mean, you may think it’s a long way down the road to the chemist’s, but that’s just peanuts to space.”

Morning Song: Someone New by Hozier.

Morning Routine: Stretching.

Breakfast: Oatmeal with Brown Sugar and Blueberries.

EVA #7: Collect. More. Rocks.

But none of that matters today. The routine, the unbelievably scenic views, the amazing science being done during this mission. None of it… Because today, after eight days… Eight long, dry shampoo filled, wipey scrubbed, hair stiff as uncooked spaghetti noodles days. I. Finally. Took. A. Shower.

While on Mars, water is a limited resource just like anything else. It’s hard to find on Mars; therefore, it must be conserved. Our mission instructions allot us a 2 minute Navy shower every 2-3 days. The instructions for a 2 minute Navy shower: Step 1: Turn on the water. Step 2: Jump in the water. Step 3: Rinse and jump out. Step 4: Turn off the water. Perform Steps 1-4 in under 45 seconds. Step 5: Scrub your body with soap in 15 to 30 seconds. Step 6: Repeat Steps 1-4 to rinse the soap off. It is a stark change from my 45 minute showers back on Earth, and it is not particularly pleasant, but it gets the job done.

But why did I wait 8 days? Our habitat holds enough water for 1-2 people to take a shower per day. Added bonus, our water heater has a mean personality and rarely works, no matter how nicely we ask. Our Engineer caved first. She described her shower has having a brief moment of warmth. Our Commander and Executive Officer experienced the same sensations of freezing water, pleasant warmth, and scalding, burn your skin off, temperature, then back to freezing cold. Our poor Geologist only felt the ice chattering cold. There was one common reaction to taking a shower across the crew, we have been reborn. It is as if you are a new person, slipping into a fresh skin like a recently molted lizard. Nobody particularly smells on this crew (or maybe we have just gotten used to each other’s smells), but in an effort to conserve water, I endured the dry shampoo and baby wipe scrub downs until I couldn’t possibly take it anymore! And I swear my hair has never felt so soft as it does today. In fresh clean clothes. With my fresh clean body.

That’s it for today. That was the highlight. We are always thinking of our loved ones. We hope they don’t miss us too much, and we want them to know that we are having a fantastic time on the Red Planet. We can’t wait to share all of our adventures with them when we return… And take a 45 minute shower. With working warm water. And ice cream.

Movie (or Show) Answer: Hitchhiker’s Guide to the Galaxy

Operations Report – January 6th

SOL: 8

Name of person filing report: Kasey Hilton

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator (hours run): 17hr 58min; Turned on last night (05Jan2019) at 15:05; Turned off this morning (06Jan2019) at 9:03; Turned on tonight (06Jan2019) at 13:24

Solar SOC – Turned on (05Jan2019) 80%; Turned off (06Jan2019) 80%; Turned on (06Jan2019) 72%

Diesel Reading – 73%

Propane Reading – 43%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 70%; 345 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 70%; 215 gallons (assuming that the tank is 300 gallons)

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 19:33 to refill tank

Water Meter: 01398548

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes

Hours: 65.5

Beginning charge: 100%

Ending charge: 55%

Currently charging: Yes

Opportunity rover used: No; Still out of brake fluid

Hours: 45.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 63.7

Beginning charge: 100%

Ending charge: 71%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: Not applicable

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: All ATVs were started and run to prevent the batteries from freezing

HabCar used and why, where? Not used

CrewCar used and why, where? Off site

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Batteries in suits #7 and #8 were replaced since each of their batteries had a voltage less than 12V

Summary of Hab operations: EVA airlock light fixed

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

GreenHab Report – January 6th

Greenhab Officer:

Jake Qiu

Environmental Control:

Heating

Cooling w/ ambient air (1hrs)

40% Shade Cloth on

80% Shade Cloth on

Average Temperatures:

Low: 23.1°C

High: 37.6°C

Hours of Supplemental Light:

5

Daily Water Usage of Crops:

7 gallons

Water in Blue Tank (lbs):

74%

Times of Watering for Crops:

0936

1529

Changes to crops:

* Cucumbers starting to look worse

* Sunflowers look ok but some of the leaves are starting to shrivel up.

* 1 pot of the spinach is starting to look worse

* Tomatoes looks the same

* Other plants showing progress

Narrative:

Watered plants throughout the day as needed. Some of the pots were still damp but dryer

in comparison to yesterday. Light watering was done. I am thinking about adding fertilizer (miracle grow)

to plants that look worse.

Experiment Results:

Date: SOL8 Crew 202 (SOL14 overall)

Watered all 21 trays as needed

Watered Crew 201 experiments

Harvest:

Support/Supplies Needed:

Astronomy Report – January 5th

Crew 202 Astronomy Report 05-Jan-2019
Crew Astronomer: Cesare Guariniello

MDRS ROBOTIC OBSERVATORY
Robotic Telescope: MDRS-WF
Objects Viewed: Sending photo of Horsehead Nebula observed on 04Jan2019

Problems Encountered: none

Here is the Leo Triplet (M65 bottom right, M66 top right, NGC 3628 left). Unfortunately I caught an airplane during the exposure with the red filter.

Cesare

Commander, MDRS 202

Operations Report – January 5th

Howdy Mission Support,

Crew 202 Operations Report 05-JAN-2019

SOL: 7

Name of person filing report: Kasey Hilton

Non-nominal systems: EVA airlock light

Notes on non-nominal systems: EVA airlock light is out, guessing the light needs to be changed

Generator (hours run): 18hr 5min; Turned on last night (04Jan2019) at 15:02; Turned off this morning (05Jan2019) at 9:07; Turned on tonight (05Jan2019) at 15:05

Solar SOC – Turned on (04Jan2019) 81%; Turned off (05Jan2019) 100%; Turned on (05Jan2019) 80%

Diesel Reading – 73%

Propane Reading – 45%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 75%; 375 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 75%; 225 gallons (assuming that the tank is 300 gallons)

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 19:29 to refill tank

Water Meter: 01398288 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes

Hours: 65.1

Beginning charge: 100%

Ending charge: 45%

Currently charging: Yes

Opportunity rover used: No; Still out of brake fluid

Hours: 45.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 63.3

Beginning charge: 100%

Ending charge: 59%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: Not applicable

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: All ATVs were started and run to prevent the batteries from freezing

HabCar used and why, where? Not used

CrewCar used and why, where? Off site

General notes and comments: The trash has not been taken out since 02Jan2019 and is starting to build up a bit in the engineering airlock

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: EVA airlock light is out, I am planning on replacing the bulb in the morning; No change in power system issues from earlier this week, still having to manually start the generator and the SOC is not increasing

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Best,

Kasey

Journalist Report – January 5th

MDRS Crew 202, Journalist Report

Sol 7 – 01/05/2018

Name the space movie (or show) given the following quote. Answer at the end of the Report:

I know I’ve made some very poor decisions recently, but I can give you my complete assurance that my work will be back to normal. I’ve still got the greatest enthusiasm and confidence in the mission. And I want to help you. (Hal 9000)

A folky electronic song by Young the Giant started today’s morning along with meditation and chocolate chip pancakes. Dreams of ice cream are not going away anytime soon, but we’re going to add sweet to everything in a hopeless effort to curb the cravings.

Our 6th Extra-Vehicular Activity (EVA) was crewed by our Commander, Geologist, Engineer, and Executive Officer. Their mission: Find more rocks (there is a theme through these EVAs). The original plan called for a longer EVA to the East in an area where the canyons open up in search of “fluvial” samples, or rocks often found near former riverbeds or water. Unfortunately, winter has not been kind to Crew 202 and the snow covered most of the trail. No matter. On Mars, you must be adaptable and expect the unexpected. In lieu of the canyon view, the crew stopped along three waypoints on the visible road. The first, a large plain with several short and stout hills. A fairly standard view here on the Red Planet. The second stop was at the base of Mt. Nutella. The mountain has no known affiliation with the delightful hazelnut spread, but it is a good place to collect sand. From our south facing window in the habitat, there is a tall mountain peak in the distance poking out between two hills. While on the trail between Stop 1 and Stop 2, the entire mountain range was in full view. A beautiful sight to behold as the Crew attempted to navigate to Stop 3. Attempted being the keyword here. Have you ever used Google Maps and no matter how many times it says “recalculating”, you still can’t find the location? This was the case with Stop 3. The stop was not a complete loss as our Commander tested various risk scenarios on the crew and the Executive Officer took radioactive measurements over the area. Our Executive Officer is originally from Ukraine and an Army drill sergeant back in the U.S.A. It takes a lot to shake this man. The only comment he made towards this EVA was, “This was the coldest EVA so far”… That means it was absolutely Artic out there. I can’t say I envied my fellow crew members, but it does sound like they found great samples for our geological studies and witnessed beautiful views.

Back at the habitat, our Health & Safety Officer has been hard at work on his own research projects! Jake Qiu is our Health & Safety Officer and our Green Habitat Officer in Crew 202. In short, he keeps the people, and the plants, healthy. Jake is a biological engineering student at Purdue University back on Earth and is working on a project which analyzes how the microbiomes of people affect the microbiomes of the sterilized plants we plan to bring to future Mars missions. A microbiome is a community of bacteria that thrive on both living and non-living objects. When we talk about microbiomes on people, this could refer to the community of bacteria on your skin, in your gut, or various other places on the human body. You are COVERED in “cities” of bacteria. If your body was a bacteria population map, your belly button would be Los Angeles. When we send plants to space, we “sterilize” the plants meaning we kill any existing bacteria living on the plants. Unfortunately, we cannot completely sterilize the humans, so our microbiomes travel with us to space. Jake is growing sterilized microgreens using hydroponics to determine how the crew’s microbiomes will affect the microbiomes of the microgreens. Pretty neat right? Hydroponics is simply a plant growing system that does not use soil. This method is also used on Earth and is a popular system for growing plants inside buildings. Our method uses two gardening trays stacked on top of each other. The plant seeds are placed on top of a cloth made from wicking fabric and laid across the top gardening tray. Water is poured into the bottom gardening tray, absorbed through the cloth, and Wala! Microgreens are sprouted! Microgreens are not only great for this project since they grow quickly, they are also great for Mars! They provide essential vitamins and nutrients that are difficult to find in the shelf-stable “food” that will be the main feature of most Martian cuisine. Additionally, Jake intends to answer whether the interaction of our microbiomes will affect the physical being of the plants (i.e. the size of the plant leaves)? Are there pathogens, or diseases, present in the microgreen microbiomes and could that transfer to humans? This is just one of the many incredible research projects in work by Crew 202. Keep up the great work Jake!

As a final note, Martians should definitely consider bringing Red Lobster Cheddar Biscuits. Shelf stable and a wonderful change of pace from the soup heavy Mars diet!

Movie (or Show) Answer: 2001: A Space Odyssey

Greenhab Report – January 5th

Hi Mission Support,

Crew 202 Greenhab Report 05-Jan-2019
Greenhab Officer:
Jake Qiu

Environmental Control:
Heating
Cooling w/ ambient air (1hrs)

40% Shade Cloth on
80% Shade Cloth on

Average Temperatures:
Low: 25.1°C
High: 35.0°C

Hours of Supplemental Light:
5

Daily Water Usage of Crops:
6 gallons

Water in Blue Tank (lbs):
78%

Times of Watering for Crops:
0736
1627

Changes to crops:
* Cucumber leaves did not perk back up early in the morning before sunrise (0611)
* 2 pots of spinach lacking progress
* 1 snap pea pot lacking progress
* Majority of other plants displays good progress

Narrative:
Watered plants throughout the day as needed. 3 gallons were used for microgreens and 2
gallons were used for other plants. Was not sunny today and there was moisture still
visible in some pots which I did not water. For cucumbers and tomatoes, I did not water.
Will determine if more water is required tommorow.

Experiment Results:
Date: SOL7 Crew 202 (SOL14 overall)

Watered all 21 trays as needed

Harvest:

Support/Supplies Needed:

Best Regards,

Jake

EVA Report – January 5th

Crew 202 EVA Report 5-Jan-2019

EVA #6

Author: Denys Bulikhov (EXO)

Purpose of EVA: Collection of geological samples and ambient radiation readings

Start time: 11:03

End time: 13:13

Narrative: EVA 6 crew took Cow Dung Road up to the Galileo Rd and then down to Cactus road. Crew stopped in the middle of the valley where they collected geological samples and ambient radiation readings. Due to snow, collection of samples was problematic, and further movement down the Cactus road was impossible. EVA 6 was cut short. On the way back, crew stopped at Mtn. Nutella to take pictures. After that they returned to the Hab.

Destination: Cactus Rd.

Coordinates: N520500, E4252000

Participants: Cesare Guariniello (CMD), Denys Bulikhov (EXO), Ellen Czaplinski (GEO), Kasey Hilton (ENG)

Road(s) and routes per MDRS Map: Cow Dung Road, Galileo Rd, Cactus Rd.

Mode of travel: Driving and walking

Vehicles used: Spirit and Curiosity.

With best regards,

Denys Bulikhov